OpenStudy (anonymous):

2/3(6x-9)=-2+3x How do i solve this?!

OpenStudy (anonymous):

I would like to know also

hero (hero):

I would show you but you don't like my methods apparently

hero (hero):

I am grown up, I'm just stating the facts.

hero (hero):

But let me just show you this one thing and if you don't like it, you can just ask someone else.

hero (hero):

So you have: \(\frac{2(6x-9}{3} = -2+ 3x\)

hero (hero):

\(\frac{2(6x-9)}{3} = -2 + 3x\)

OpenStudy (anonymous):

I thought it was (2/3)(6x-9)

OpenStudy (anonymous):

^what zarby said...

hero (hero):

The first time you posted the fractions, you posted them linearly and posted them the exact same way but the terms were supposed to be in the numerator. To eliminate the confusion in inconsistency, use the drawing tool or learn latex.

OpenStudy (anonymous):